Standard deviation versus coefficient of variation as measures of risk Greengage, Inc., a successful nursery, is consideringPlease assist in solving all parts
Answer #1

a) Project B is the least risky project as the range is the minimum. Range is a measure of dispersion. More the range, more risky is the project, and vice versa.

b) Project A has the lowest standard deviation. Standard measure is not an appropriate measure of risk as it is an absolute measure of dispersion. Risk should be compared together with returns of the project so as to know risk per unit of expected return.

c) Coefficient of Variation = (Standard Deviation/ Expected Returns) x 100

Project A - C.V = (3.5 / 12.7) x 100 = 27.56%

Project B - C.V = (3.9 / 13.3) x 100 = 29.32%

Project C - C.V = (3.8 / 11.7) x 100 = 32.48%

Project D - C.V = ( 3.7/ 11.7) x 100 = 31.62%

Greengage's owner should select Project A as the Coefficient of Variation is the minimum, which means the relative risk, when compared together with expected returns, is the minimum.